# Can one size fit all?: Measuring Failure in Multi-Document Summarization Domain Transfer

### v1

@@ Abstract
Abstractive multi-document summarization (MDS) is the task of automatically summarizing information in multiple documents, from news articles to conversations with multiple speakers. The training approaches for current MDS models can be grouped into four approaches: end-to-end with special pre-training (“direct”), chunk-then-summarize, extract-then-summarize, and inference with GPT-style models. In this work, we evaluate MDS models across training approaches, domains, and dimensions (reference similarity, quality, and factuality), to analyze how and why models trained on one domain can fail to summarize documents from another (News, Science, and Conversation) in the zero-shot domain transfer setting. We define domain-transfer “failure” as a decrease in factuality, higher deviation from the target, and a general decrease in summary quality. In addition to exploring domain transfer for MDS models, we examine potential issues with applying popular summarization metrics out-of-the-box.
@@ 1 Introduction
Figure 1: Sample generations from SummN trained on Multi-XScience (SummN-Science) and PRIMERA trained on Multi-News (PRIMERA-News) on the ConvoSumm-Reddit dataset (Conversation). Relevance is measured automatically with UniEval.
@@
Recent MDS system surveys compare different models either on a single domain (e.g., conversation/dialogue (Zhang et al., 2021; Mullick et al., 2024)) or focus on MDS corpora biases (Wolhandler et al., 2022; Dey et al., 2020). While Zhang et al. (2021) found that retrieve-then-summarize models are best according to ROUGE, Mullick et al. (2024) did not find a clear pattern in model performance across a variety of metrics. Few works explore the zero-shot domain transfer ability of MDS systems, but Zhang et al. (2021) did find that pre-training on news datasets can transfer well to conversation. However, Fabbri et al. (2021a) a zero-shot model from news to conversation had lower performance than on conversation alone.
An example is in Figure 1, where two models, one trained on Multi-News (Fabbri et al., 2019) and the other on Multi-XScience (Lu et al., 2020) are evaluated on a sample from ConvoSumm-Reddit (Fabbri et al., 2021a). The difference in writing style is immediately apparent: the news-trained model draws out quotes, the science-trained model inserts citations, and both are completely different from the expected style of the Reddit conversation summary. The difference in style is also captured by similarity metrics ROUGE (Lin, 2004) and UniEval-Relevance (Zhong et al., 2022)–while the ROUGE scores, which capture word overlap, only differ by a few points, the semantic-based, trained Relevance scores vary widely.
@@
Figure 1 highlights the three axes of domain transfer for MDS systems: 1) domain of the training data, 2) type of model, and 3) evaluation metrics. While there are popular benchmarks for summarization, including the notable SummEval (Fabbri et al., 2021b), few are for multi-document summarization, and most datasets focus on the news domain. We choose three high-quality human-written datasets to represent News (Multi-News (Fabbri et al., 2019)), Science (Multi-XScience (Lu et al., 2020)), and Conversation (ConvoSumm-Reddit (Fabbri et al., 2021a)) domains.
@@
  • Smaller language models (BART-size) hold more stylistic bias with their training data;
  • Within News, Science, and Conversation, there is no clear “best” domain for general transfer;
  • Lexical and semantic similarity metrics both fail to capture summary quality;
  • Current non-reference-based metrics do not scale well to the multi-document setting;
  • Trained evaluators have bias and penalize out-of-domain datasets;
Figure 2: Summary of all the domains (datasets), multi-document summarization (MDS) model types, and evaluation metrics analyzed in this work.
@@ 2 Domain Transfer of Multi-Document Summarization
While there is in-depth work on analyzing summarization and MDS systems and corpora (Zhang et al., 2021; Mullick et al., 2024; Fabbri et al., 2021b; Dey et al., 2020), there is no systematic evaluation of how each model type performs when applied out-of-domain and how the automatic evaluation metrics can over- or under-represent model performance. We draw from existing models and tools to explore the intricacies of domain transfer with MDS models and their evaluation methods. The main dimensions of our analyses are domain (i.e., dataset), model type, and evaluation metrics. See Figure 2 for summary.
@@ 2.1 Summarization Dataset Domains
Multi-News (Fabbri et al., 2019) is a news-based MDS dataset featuring 56,216 human-written summaries from Newser.222newser.com
@@
2.2 Multi-Document Summarization Model Types
Current high-performing fine-tuned abstractive summarization methods can be grouped into four approaches: summary-specific pre-training (Zhang et al., 2020a; Xiao et al., 2022a; Peper et al., 2024), incremental summarization (“chunk-then-summarize”) (Zhang et al., 2022), extracting claims or viewpoints (“extract-then-summarize”) (Fabbri et al., 2021a; Ernst et al., 2022; Ouyang et al., 2023), and fine-tuning or few-shot summarization with large pre-trained models (Mullick et al., 2024). These groupings have been used in prior work (Mullick et al., 2024).
@@ 2.2 Multi-Document Summarization Model Types
Prior work has shown that fine-tuning a language model specifically pre-trained with summary-focused objectives improves downstream performance over non-specialized models (Xiao et al., 2022b; Zhang et al., 2020a; Peper et al., 2024). Zhang et al. (2020a) first introduced this concept of summary-specific pre-training, and it was later augmented to the multi-document case (Xiao et al., 2022b; Peper et al., 2024). We select PRIMERA to represent this group. PRIMERA is Longformer-Encoder-Decoder model initialized from BART-Large (Beltagy et al., 2020) and pre-trained with an Entity Pyramid objective on a large news corpora (Xiao et al., 2022b). While PELMS (Peper et al., 2024) adjusts the pre-training strategy to improve upon PRIMERA, we omit comparison to this model because it also uses a data filtering step, which overlaps with the extract-then-summarize approach.
@@ 2.3 Summarization Evaluation
Other metrics in the spirit of ROUGE/BLEU are semantic-based instead of lexical, which do not penalize generations with similar meanings but different word choices. The most popular version is BERTScore (Zhang et al., 2020b), which maps tokens from the generation to the reference before calculating overlap. An alternative is MoverScore, which “optimized” version of BERTScore (Zhang et al., 2020b) which uses soft token alignments from the hypothesis to the reference instead of a hard 1:1 mapping, allowing for a greater reward for semantically similar generated text.
Other similarity metrics are sub-“dimensions” of a trained multi-dimensional evaluator, like UniEval. UniEval is a T5-Large model fine-tuned on synthetically augmented datasets for positive and negative examples of different “dimensions” of summary attributes (fluency, relevance, coherence, consistency) (Zhong et al., 2022). The UniEval-Relevance dimension is a similarity metric that compares the reference and generation as a whole instead of per token.
@@
Not all desired qualities of a summary depend on a reference, some are intrinsic or rely on the source documents (e.g., fluency). UniEval measures both fluency and coherence. UniEval-Fluency is based solely on the generation, and UniEval-Coherence measures the coherence of the generation with respect to the source.
@@
A growing concern with generated text is the possibility of hallucinations and non-faithfulness to the source material (Min et al., 2023). While objective factuality is important for the safety of LLMs (e.g., knowing that boiling water is hot), for MDS we typically assume that the source documents are factually correct and that a summary is only factual if its claims match the claims from the source. We measure summary factuality with UniEval-Consistency, a single-dimensional evaluator trained to measure the factual consistency of generated text to the source (as opposed to continuously trained as with UniEval-Fluency and UniEval-Coherence). A popular framework for factuality analysis is decompose-then-verify (Min et al., 2023; Jiang et al., 2024), and we leave this analysis to future work.
We include ROUGE, MoverScore, and multiple dimensions of UniEval in our analysis as representations for similarity to ground truth, factuality, and overall quality metrics. While BARTScore (Yuan et al., 2021) measures many of the same dimensions as UniEval, we chose UniEval because it is better correlated with human evaluation (Zhong et al., 2022) and the potential bias of BARTScore to model generations, since most of the models are BART-based (“self bias”) (He et al., 2023). Another evaluation for summaries is source coverage, or how much of the source documents are represented in the summaries (Wolhandler et al., 2022). We leave this evaluation to future work.
@@ 3 Experimental Setup
Our goal is to explore the intricacies of domain transfer with multi-document summarization models and their evaluation methods. The main dimensions of our analyses are different datasets (i.e., domains), models, and evaluation metrics (Figure 2). We chose 3 dataset domains (News, Science, Conversation), 4 model types (end-to-end/direct), chunk-then-summarize, extract-then-summarize, and pre-trained LLMs), and 7 metrics for a total of 84 points of analysis (Section 2).
@@ 3.1 Model Training
The hyperparameters for each model and dataset combination are in Appendix B. Each model was trained on each dataset, and we refer to them as ModelName-Domain, e.g., ConvoSumm fine-tuned on Multi-News is ConvoSumm-News.
@@
As input, the source documents are delineated with a special <doc-sep> token, and the model’s global attention is set on those tokens. To handle examples with source documents that are too long for the context window, we follow the author’s code and evenly truncate the documents to fit. We used the authors’ released models of PRIMERA fine-tuned on Multi-XScience and MultiNews and trained PRIMERA-ConvoSumm-Reddit with the released author code.
@@
We use a 3 course-stage process for all datasets. A separate model is trained for each coarse stage, leading to $N+1$, or $4$ fine-tuned BART models for each training dataset. An important note is due to the setup of summarizing chunks, there are no too-long context issues as with direct models like PRIMERA and all source documents are considered by the model.
@@
Unlike the other models, we do not fine-tune Llama 3.1 8B-Instruct. [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct To compare to the domain-transfer experiments, we provide an example from the training set to the model, i.e., for News-Convo, an example from Multi-News is provided to Llama and then the model is prompted to complete an example from ConvoSumm. We also include a zero-shot version which does not provide a complete example to Llama.
@@ 3.2 Evaluation Details
ROUGE scores have been known to vary depending on the implementation used. We use the HuggingFace wrapper for the Google Research implementation. [^4] https://huggingface.co/spaces/evaluate-metric/rouge We report ROUGE-1 (R1) and ROUGE-L-Sum (RL-Sum) to follow prior work. For the other evaluation metrics, MoverScore and multiple dimensions from UniEval, we use the trained models as they are without modification.
Factuality Ground-Truth Similarity Quality
Cons. (UE) Rel. (UE) R1 RLSum MoverScore Fluency (UE) Coh. (UE)
Conversation 0.68 0.75 0.30 0.20 0.07 0.91 0.72
Avg. $\Delta$ +18% -13% -9% -19% -31% -1% +4%
Ref. 0.71 0.99 1.00 1.00 1.00 0.90 0.73
News 0.72 0.78 0.43 0.22 0.20 0.91 0.79
Avg. $\Delta$7% +2% -51% -41% -92% -2% +1%
Ref. 0.72 0.99 1.00 1.00 1.00 0.92 0.80
Science 0.78 0.63 0.24 0.15 0.03 0.78 0.64
Avg. $\Delta$7% -12% +6% +1% -7% -1% +4%
Ref. 0.55 0.99 1.00 1.00 1.00 0.66 0.43
Table 1: Multi-document summarization (MDS) model performance and relative performance aggregated by domain when trained and tested on the same domain (top) and different domains ($\Delta$), respectively. Results are aggregated across all the test splits of each dataset. A positive $\Delta$ shows a performance gain and a negative $\Delta$ is a performance loss as compared to the model’s scores on the in-domain data. A relative change of over $1$ is possible with MoverScore because of its $[-1,1]$ range. Detailed results are in Table 7. “Ref” refers to the results when running the metrics on the ground-truth data.

### v2

@@ Abstract
Abstractive multi-document summarization (MDS) is the task of automatically summarizing information in multiple documents, from news articles to conversations with multiple speakers. The training approaches for current MDS models can be grouped into four approaches: end-to-end with special pre-training (“direct”), chunk-then-summarize, extract-then-summarize, and inference with GPT-style models. In this work, we evaluate MDS models across training approaches, domains, and dimensions (reference similarity, quality, and factuality), to analyze how and why models trained on one domain can fail to summarize documents from another (News, Science, and Conversation) in the zero-shot domain transfer setting. We define domain-transfer “failure” as a decrease in factuality, higher deviation from the target, and a general decrease in summary quality. In addition to exploring domain transfer for MDS models, we critically examine the behavior of various automatic evaluation metrics, including their inter-correlations and the performance of LLM-as-a-judge approaches, in these transfer settings.
Can one size fit all?: Measuring Failure in Multi-Document Summarization Domain Transfer
Alexandra DeLucia  and Mark Dredze Center for Language and Speech Processing Johns Hopkins University {aadelucia,mdredze}@jhu.edu
@@ 1 Introduction
Domain-Transfer Example
Ground-Truth A couple of the responses to this post are either satirical or sarcastic. ...
News-based Might you be a little nervous about playing a video game for free? you’re not alone. ...
Science-based @cite @cite have both mentioned that the reason for the lack of a sub is due to Microsoft’s refusal to allow sub-free games on Xbox. ...
Figure 1: Excerpt generations two types of multi-document summarization models (SummN and ConvoSumm), trained on different datasets (Multi-News+ and Multi-XScience), and both evaluated on a conversation summarization task (ConvoSumm-Reddit). The ROUGE-LSum of the generations to the Ground-Truth are 0.13 and 0.12, respectively, and UniEval-Relevancy scores of 0.78 and 0.3, respectively.
Recent MDS system surveys compare different models either on a single domain (e.g., conversation/dialogue (Zhang et al., 2021; Mullick et al., 2024) or focus on MDS corpora biases (Wolhandler et al., 2022a; Dey et al., 2020). While Zhang et al. (2021) found that retrieve-then-summarize models are best according to ROUGE, Mullick et al. (2024) did not find a clear pattern in model performance across a variety of metrics. Few works explore the zero-shot domain transfer ability of MDS systems, but Zhang et al. (2021) did find that pre-training on news datasets can transfer well to conversation. However, Fabbri et al. (2021a) a zero-shot model from news to conversation had lower performance than on conversation alone.
An example is in Figure˜1, where two models, one trained on Multi-News+ (Choi et al., 2024; Fabbri et al., 2019) and the other on Multi-XScience (Lu et al., 2020) are evaluated on a sample from ConvoSumm-Reddit (Fabbri et al., 2021a). The difference in writing style is immediately apparent: the news-trained model has a blogger-like style, the science-trained model inserts citations, and both are completely different from the expected style of the Reddit conversation summary. The difference in style is also captured by similarity metrics ROUGE (Lin, 2004) and UniEval-Relevance (Zhong et al., 2022)–while the ROUGE scores, which capture word overlap, only differ by a few points, the semantic-based, trained Relevance scores vary widely.
@@
Figure˜1 highlights the three axes of domain transfer for MDS systems: 1) domain of the training data, 2) type of model, and 3) evaluation metrics. While there are popular benchmarks for summarization, including the notable SummEval (Fabbri et al., 2021b), few are for multi-document summarization, and most datasets focus on the news domain. We choose three high-quality human-written datasets to represent News (Multi-News (Fabbri et al., 2019)), Science (Multi-XScience (Lu et al., 2020)), and Conversation (ConvoSumm-Reddit (Fabbri et al., 2021a)) domains.
@@
  • LLMs (e.g., Llama 3.1) demonstrate robust quality and factuality in domain transfer with minimal performance change, and their outputs are less influenced by few-shot examples compared to the stronger stylistic bias observed in smaller, fine-tuned models;
  • While no single source domain (News, Science, or Conversation) proves universally superior for general transfer, the stylistic nature of summaries learned during training significantly impacts cross-domain performance;
  • Lexical similarity metrics (e.g., ROUGE, MoverScore) show weak correlation with semantic relevance (e.g., UniEval-Relevance) and other quality dimensions;
  • Source-dependent non-reference metrics face scaling challenges moving from single-document summarization to multi-;
  • Trained multi-dimensional evaluators exhibit intra- and inter-correlations with each other and within dimensions;
  • The number of source documents surprisingly shows weak and inconsistent correlations with factuality and coherence metrics in domain-transfer settings across different model architectures;
Figure 2: Summary of all the domains (datasets), multi-document summarization (MDS) model types, and evaluation metrics analyzed in this work. The “Relevance” metric is in both Ground-Truth Similarity and Overall Quality groups because the UniEval (UE) and Prometheus (P) definitions are different (see Figures˜10 and 11 for details).
@@ 2 Domain Transfer of Multi-Document Summarization
While there is in-depth work on analyzing summarization and MDS systems and corpora (Zhang et al., 2021; Mullick et al., 2024; Fabbri et al., 2021b; Dey et al., 2020), there is no systematic evaluation of how each model type performs when applied out-of-domain and how the automatic evaluation metrics can over- or under-represent model performance. We draw from existing models and tools to explore the intricacies of domain transfer with MDS models and their evaluation methods. The main dimensions of our analyses are domain (i.e., dataset), model type, and evaluation metrics. See Figure˜2 for summary.
@@ 2.1 Summarization Dataset Domains
Multi-News+ (Choi et al., 2024), an automatically cleaned version of Multi-News (Fabbri et al., 2019), is a news-based MDS dataset featuring human-written summaries from Newser.222newser.com. Multi-News+ contains 44,668/5,585/5,584 entries in the train/validation/test sets.
@@
2.2 Multi-Document Summarization Models
Current high-performing fine-tuned abstractive summarization methods can be grouped into four approaches: summary-specific pre-training (Zhang et al., 2020a; Xiao et al., 2022; Peper et al., 2024), incremental summarization (“chunk-then-summarize”) (Zhang et al., 2022), extracting claims or viewpoints (“extract-then-summarize”) (Fabbri et al., 2021a; Ernst et al., 2022; Ouyang et al., 2023), and fine-tuning or few-shot summarization with large pre-trained models (Mullick et al., 2024). These groupings have been used in prior work (Mullick et al., 2024).
@@ 2.2 Multi-Document Summarization Models
Prior work has shown that fine-tuning a language model specifically pre-trained with summary-focused objectives improves downstream performance over non-specialized models (Xiao et al., 2022; Zhang et al., 2020a; Peper et al., 2024). Zhang et al. (2020a) first introduced this concept of summary-specific pre-training, and it was later augmented to the multi-document case (Xiao et al., 2022; Peper et al., 2024). We select PRIMERA to represent this group. PRIMERA is Longformer-Encoder-Decoder model initialized from BART-Large (Beltagy et al., 2020) and pre-trained with an Entity Pyramid objective on a large news corpora (Xiao et al., 2022). While PELMS (Peper et al., 2024) adjusts the pre-training strategy to improve upon PRIMERA, we omit comparison to this model because it also uses a data filtering step, which overlaps with the extract-then-summarize approach.
@@ 2.3 Summarization Evaluation
Other metrics in the spirit of ROUGE/BLEU are semantic-based instead of lexical, which are more robust to passages with similar meanings but different word choices. The most popular version is BERTScore (Zhang et al., 2020b), which maps tokens from the generation to the reference before calculating overlap. An alternative is MoverScore, an “optimized” version of BERTScore which uses soft token alignments from the hypothesis to the reference instead of a hard 1:1 mapping, allowing for a greater reward for semantically similar generated text.
Other similarity metrics are sub-“dimensions” of a trained multi-dimensional evaluator, like UniEval (Zhong et al., 2022). UniEval is a T5-Large model fine-tuned on synthetically augmented datasets for positive and negative examples of different “dimensions” of summary attributes (fluency, relevance, coherence, consistency) (Zhong et al., 2022). The UniEval-Relevance dimension is a similarity metric that compares the reference and generation as a whole instead of per token.
@@
Not all desired summary qualities depend on a reference; some are intrinsic or rely on the source documents. Intrinsic metrics include fluency (i.e., well-written and grammatical), coherence (i.e., overall flow), and the summarization-specific metrics of compression ratio (i.e., length of source documents compared to length of the summary) and source coverage (i.e., how well the summary represents the source documents). UniEval measures both fluency and coherence. UniEval-Fluency is based solely on the generation, and UniEval-Coherence measures the coherence of the generation with respect to the source.
Similar to UniEval, Prometheus (Kim et al., 2024) is another multi-dimensional evaluator, but is a general-purpose feedback model that can judge any model output according to a provided rubric. In this respect, Prometheus is essentially a distilled “LLM-as-a-Judge” model (Zheng et al., 2023). For a more up-to-date automatic evaluation method, we use Prometheus to grade the summaries on the following criteria, drawn from SummEval (Fabbri et al., 2021b): Relevance, Coherence, and Fluency. An important note is that despite the same metric names, the definitions from SummEval differ slightly from those in UniEval.
To measure compression ratio, we include both Compression-Sentence, which compares the ratio of the number of sentences in the summary to those in the source documents, and Compression-Word, which compares the ratio of the number of words in the summary to those in the source documents (Koh et al., 2022). For source coverage (Wolhandler et al., 2022a), we developed a prompt for Prometheus, Prometheus-Coverage.
@@
A growing concern with generated text is the possibility of hallucinations and non-faithfulness to the source material (Min et al., 2023). While objective factuality is important for the safety of LLMs (e.g., knowing that boiling water is hot), for MDS we typically assume that the source documents are factually correct and that a summary is only factual if its claims match the claims from the source. This definition of “factually” is better described as grounding in the source text. We measure summary factuality with UniEval-Consistency, a single-dimensional evaluator trained to measure the factual consistency of generated text to the source (as opposed to continuously trained as with UniEval-Fluency and UniEval-Coherence), Prometheus-Consistency, and Prometheus-Factuality. The Prometheus-Consistency prompt is drawn from SummEval, and the Prometheus-Factuality prompt is from the Prometheus repository.
[…]
@@ 3 Experimental Setup
Our goal is to explore the intricacies of domain transfer with multi-document summarization models and their evaluation methods. The main dimensions of our analyses are different datasets (i.e., domains), models, and evaluation metrics (Figure˜2). We chose 3 dataset domains (News, Science, Conversation), 4 model types (end-to-end/direct), chunk-then-summarize, extract-then-summarize, and pre-trained LLMs), and 15 metrics for a total of 180 points of analysis (Section˜2).
@@ 3.1 Model Training
The hyperparameters for each model and dataset combination are in Appendix˜B. Each model was trained on each dataset, and we refer to them as ModelName-Domain, e.g., ConvoSumm fine-tuned on Multi-News+ is ConvoSumm-News.
@@
As input, the source documents are delineated with a special <doc-sep> token, and the model’s global attention is set on those tokens. To handle examples with source documents that are too long for the context window, we follow the author’s code and evenly truncate the documents to fit. We used the authors’ released PRIMERA-MultiXScience model and trained PRIMERA-Convo and PRIMERA-News with the released code. [^3] While there is a PRIMERA model trained on Multi-News, we trained a new one on Multi-News+.
@@
We use a 3 course-stage process for all datasets. A separate model is trained for each coarse stage, leading to $N+1$, or 444 fine-tuned BART models for each training dataset. An important note is due to the setup of summarizing chunks, there are no too-long context issues as with direct models like PRIMERA and all source documents are considered by the model.
@@
Unlike the other models, we do not fine-tune Llama 3.1 8B-Instruct. [^4] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct To compare to the domain-transfer experiments, we provide an example from the training set to the model, i.e., for News-Convo, two randomly selected examples from Multi-News+ is provided to Llama and then the model is prompted to summarize an example from ConvoSumm. We also include a zero-shot version which does not provide a complete example to Llama, only the instruction.
@@ 3.2 Evaluation Details
ROUGE scores have been known to vary depending on the implementation used. We use the HuggingFace wrapper for the Google Research implementation. [^5] https://huggingface.co/spaces/evaluate-metric/rouge We report ROUGE-1 (R1) and ROUGE-L-Sum (RL-Sum) to follow prior work. For the other evaluation metrics, MoverScore and multiple dimensions from UniEval, we use the trained models as they are without modification. For evaluation with Prometheus, we use Prometheus2-BGB-8x7B  [^6] https://huggingface.co/prometheus-eval/prometheus-bgb-8x7b-v2.0 as an absolute judge with dimension-specific rubrics we based off the definitions from SummEval (Fabbri et al., 2021b). See Appendix˜C for details. Due to the intense compute required for Prometheus, we sampled 100 items from each dataset for all Prometheus-based metrics.
ConvoSumm PRIMERA SummN Llama 3.1
Llama 3.1
(0-shot)
Avg. $\Delta$ Avg. $\Delta$ Avg. $\Delta$ Avg. $\Delta$ Avg. $\Delta$
Factuality Consistency (UE) 0.67 -2% 0.82 0% 0.65 +3% 0.86 0% 0%
Factuality (P) 2.04 -16% 1.79 +3% 2.02 -6% 3.23 0% 0%
Consistency (P) 2.14 -15% 1.89 +1% 2.08 -7% 3.40 +1% 0%
Ground-Truth Similarity Relevancy (UE) 0.66 -3% 0.48 +6% 0.66 -19% 0.92 0% 0%
ROUGE-1 0.35 -27% 0.27 +12% 0.35 -30% 0.29 +10% 0%
ROUGE-LSum 0.20 -28% 0.14 +6% 0.20 -28% 0.17 +4% +1%
MoverScore 0.10 -86% 0.06 -332% 0.10 -88% 0.10 +35% -2%
Quality Relevancy (P) 2.36 -15% 2.18 +3% 2.33 -6% 3.53 +1% +1%
Fluency (UE) 0.79 -5% 0.69 +2% 0.76 0% 0.94 0% 0%
Fluency (P) 2.30 -21% 1.67 +9% 2.26 -5% 3.97 0% 0%
Coherence (UE) 0.60 -14% 0.70 +2% 0.63 -2% 0.95 0% 0%
Coherence (P) 2.31 -20% 1.63 +8% 2.31 -11% 3.95 0% 0%
Comp.-Sent 0.71 +7% 0.66 +1% 0.69 +9% 0.64 +4% 0%
Comp.-Word 0.74 +1% 0.55 +2% 0.70 -1% 0.55 +2% 0%
Coverage (P) 1.93 -12% 1.83 +3% 1.95 -6% 3.03 +1% 0%
Table 1: Multi-document summarization (MDS) model performance and relative performance aggregated by model when trained and tested on the same domain (left) and different domains (right, $\Delta$), respectively. A relative change of over 111 is possible with MoverScore because of its [1,1][-1,1][ - 1 , 1 ] range. Llama (0-shot) is compared to Llama with in-domain examples.